3 Egyptian band, West el Balad, have just finished recordin their latest CD. Their record companys marketing departmernt d
Answer #1
P Q TR=P*Q MR
24 10000 240000
22 20000 440000 20
20 30000 600000 16
18 40000 720000 12
16 50000 800000 8
14 60000 840000 4

the total revenue is calculated by multiplying the price of the commodity with quantity demanded

TR = Price*Quantity (P*Q). The marginal revenues are calculated by the formula

MR = Change in TR/Change in Q

for MR corresponding to demand of 20000 units, MR = (440000-240000)/(20000-10000) = 20